A stock has an expected return of 8 percent, its beta is 0.5, and the risk-free rate is 3.6 percent. The expected return on the market must be Question 2 options: 10.2% 14.5% 15.2% 11.1% 12.4%
Answer #1

expected return=risk free rate+beta*(market rate-risk free rate)

8=3.6+0.5*(market rate-3.6)

(8-3.6)=0.5*(market rate-3.6)

market rate=(8-3.6)/0.5+3.6

=12.4%